[figure: 9.7 cm]

EFFECT OF PRESSURE ON MAGNETIC PHASE TRANSITIONS AND RELATED VOLUME ANOMALIES OF

Ce_{2}Fe_{17} COMPOUND

Z. Arnold^{a}, I.V. Medvedeva^{b}, O. Prokhnenko^{a}, J. Kamarád^{a}, A.G. Kuchin^{b},

aInstitute of Physics AS CR, Cukrovarnická 10, 162 53 Praha 6, Czech Republic

bInstitute of Metal Physics, Ural Division of Russian Academy of Sciences, S.Kovalevskaya 18, 620219 Ekaterinburg, Russia.

To get a better insight into the volume dependence of the magnetic properties on Ce_{2}Fe_{17} compound we performed a.c. susceptibility and thermal expansion studies in temperature range 10 – 320 K and in pressure range up to 10 kbar. At ambient pressure T_{C} = 94 K and T_{N} = 206 K. Our thermal expansion measurements at ambient pressure using the strain gage method showed that both these transitions are connected with remarkable volume anomalies, showing the spontaneous volume magnetostriction ω_{S} ≈ 1% at 10 K. The behaviour under pressure is rather complex with giant pressure effect on Curie temperature, dT_{C} /dp = −38 K/kbar, with suppression of ferromagnetic order at 2.5 kbar and moderate pressure effect on Néel temperature, dT_{N} /dp = −1.7 K/kbar. Our compressibility and thermal expansion measurements showed that the T_{C} volume anomaly is suppressed by the pressure of about 2 kbar and the total spontaneous volume magnetostriction disappears at a pressure of 4 kbar. We ascribe this behaviour to the strong sensitivity of magnetic structures of Ce_{2}Fe_{17} to interatomic distances. We acknowledge support from Projects No. 106/99/0183 GA CR and A1010811 GA AS CR and No.98-02-16165 RFBR.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr. Zdenek ARNOLD

Address for correspondence : Institute of Physics, AS CR Cukrovarnická 10 162 53 Praha 6 Czech Republic

Email address : arnold@fzu.cz